# AppSupplemental Materials References

### v1

@@ 7 Do Legislators Respond to Shootings in Other Ways?
In the California State Legislature, legislators frequently sign bills as a cosponsor as a means of signaling support for a bill in a low-cost manner. We manually collected the sponsor and cosponsor lists of all firearm-related bills in both chambers of the California State Legislature from 2011 to 2022. Following Fowler (2006), we construct a directed cosponsorship network, with each edge drawn from a cosponsor of a bill to its sponsor. Figure S.11.1 describes this network. This network is densely interconnected, particularly among Democrats, suggesting partisan homophily in the sponsorship network. Table S.11.1 records the most “important” legislators in each session and chamber, measured by highest Outdegree (the number of bills a given legislator cosponsors in a year) centrality, Betweenness centrality, and Eigenvector centrality.
How might mass shootings affect the sponsorship and cosponsorship behaviors of California lawmakers? Beyond Outdegree (which only captures the cosponsorship activities), we define “Activity” as the fraction of firearm-related bills each legislator sponsors and cosponsors out of all firearm-related bills in a year. Together, these measures serve as proxies for a legislator’s level of activity in the gun policy domain, along distinct albeit complementary dimensions. In Table S.11.2, we report the results of our staggered DiD on these two outcomes of interest — Activity and Outdegree. We only find a statistically significant decrease in both the overall Activity measure and the Outdegree measure for treated Democrats, relative to other Democrats, following a shooting in the first post-treatment period. This suggests that treated Democratic legislators in California are less active in the gun policy domain, on average, in the year immediately following a mass shooting. This estimate, however, must be taken with a grain of salt considering possible pre-trending observed in the pre-treatment periods, suggesting a violation of the parallel trends assumption.
Of course, it is possible that legislators engage in other forms of behavior outside of the legislative process that we may fail to capture in this paper. In subsequent work, we aim to consider the effects of mass shootings on the public-facing behavior of lawmakers — such as social media activity and visits to victims — rather than just the more “official” forms of behavior our current study focuses on.
@@ 8 Conclusion
In California, we find a statistically insignificant 1.45% reduction in treated legislator’s support for restrictive firearm-related policies, compared to all other legislators, an insignificant 4.8% reduction among California Democrats, and an insignificant 0.9% increase among California Republicans. This null finding persists when we expand our definition of treatment to consider the role of geographic proximity and the effects of mass shootings on the voting behavior of legislators who represent districts near the shooting’s location and not just the district in which the shooting occurred. When we expand our analysis to include Colorado, Florida, Georgia, Missouri, and Texas, we again estimate statistically insignificant effects of mass shootings on legislators compared to members of their own party. Our scoring regime ultimately covers a total of 1,975 unique, individual legislators across six states from 2011 to 2022.
@@
Our findings challenge canonical theories that posit focusing events, such as mass shootings, as critical junctures for policy change. Instead, the results highlight the limits of these events in altering entrenched partisan dynamics, suggesting that focusing events may have differential impacts based on the issue domain and political context. Legislators we interviewed told us that mass shootings led them to sponsor additional gun safety legislation, or reconsider how they could best represent the victims of mass shootings, but also admitted that pressures from party leaders constrained their autonomy in responding to the tragedy at hand. Others suggested that political differences within their constituencies led them to walk a “fine line” as a representative — they could not push a policy agenda at odds with the wishes of the families of victims, for instance. This was in stark contrast to others who told us that the tragedy offered an opportunity to push through policy changes deemed “great” by some legislators and “unconstitutional” Trojan horses by others.
@@
This nuance helps contextualize our finding that shootings in particular periods of time may produce differential effects. Some scholarship suggests that particular mass shootings may lead to differential responses, with characteristics such as the race of victims leading to disparate legislative reactions or swings in public opinion (Markarian, 2024; Walker, Collingwood and Bunyasi, 2020). Future research should disaggregate the effects of individual shootings to explore whether high-profile events produce unique responses among legislators. Additionally, future studies focusing on other policy domains could assess whether the limited effect estimated in this study is unique to firearm-related issues or generalizable to other policy areas marked by polarization. In this sense, policy-specific ideal points estimates need not be tailored to firearms-related policies alone.
@@
Partisan polarization has contributed to a stagnant firearms policy environment at the federal level, and our findings suggest that this may persist at the sub-national level both within the California Legislature and across five other states with varying gun policy environments. Our results suggest that while state legislatures are active on gun policy issues, mass shootings do not cause any measurable changes in individual-level voting behaviors of elected officials, on average. Every legislator we spoke with said, at different points, that they felt a degree of responsibility to act in the aftermath of the shooting, and yet our results suggest that these events, on average, do not cause a legislator to change their positions on firearm-related policies. This does not, however, indicate that a state’s firearm policy environment is necessarily stagnant — shootings may be changing the scope of legislation that lawmakers vote on (Luca, Malhotra and Poliquin, 2020), but that is beyond the scope of our current study.
@@
This has mixed implications for the state of representative democracy, broadly speaking. On the one hand, the stagnation we observe in the voting behavior of state legislators suggests that they are not accountable to focusing events that may otherwise be thought to incite changes in their behaviors. On the other hand, perhaps this is to be expected, given the relative stability of public opinion on firearm policies after mass shootings. Regardless, while we do not intend to claim that individual votes on niche policies by state legislators capture the attention of the mass public, the outcomes of these votes do have long-lasting and far-reaching implications for accountability and representation.
@@
S.1 Qualitative Interview Procedures
@@
S.2 Locations and Details of Mass Shootings
In this paper, we consider the effect of mass shootings on legislative behavior in California from the 2011-2012 legislative session to the 2021-2022 session. To that end, we focus on shootings occurring between 2011 and 2022 in the state, though we also include the 2010 North Hollywood shooting that may have reasonably affected legislators serving in the 2011-2012 session. As these legislators would reasonably be considered treated in these early sessions, we cannot use them as control units in our later estimations. Table S.2.1 identifies the shootings that are included in our study — specifying the date, location, number of victims (not including the shooter), and the State Senate and Assembly districts in which the shooting occurred (accounting for redistricting following the 2010 Census).
Table S.2.1: Shootings Included in Study and Districts Immediately Affected
@@
  • Note: Data on shooting location and number of victims (not including the shooter) comes from the Violence Project. Upper refers to the upper chamber district in which the shooting occurred and Lower refers to the respective lower chamber district. Both of these districts are identified through mapping the latitude/longitude location of a shooting to the appropriate state legislative district shapefiles from the United States Census Bureau. These locations are cross-referenced with records from the Gun Violence Archive, when available. The 2010 North Hollywood, CA shooting is effectively dropped from our analyses – as we do not have pre-shooting observations, we cannot consider the effects of the shooting on the voting behavior of affected legislators, so the legislators representing this district that appear in our data from 2011-2022 are considered as previously treated, and removed from subsequent models. Post-Treat reports the year that is considered as $t=0$ in our formal models, and is assigned as the year of the shooting date if the date of the shooting occurs within the first six months of the year, or the subsequent year following the year of the shooting date if the shooting occurs within the last six months of the year. For shootings occurring before the redistricting done following the 2010 Census, we report the district that was affected under the earlier maps, and cross-reference district boundaries after the Census to ensure that legislators who were redistricted to represent a new area are still considered treated if they previously represented an affected district. District shapefiles from this earlier period are sourced primarily from the Redistricting Data Hub.
@@
S.3 Validating the GLC Legislative Scoring Regime
@@
Figure S.3.1: Number of Firearm-Related Bills Introduced in California Over Time, by Effect Type
Note: Counts of bills by original GLC coding, with missing observations corrected according to RA scoring.
Figure S.3.2: Number of Firearm-Related Bills Introduced in All States in Study, Over Time, by Effect Type
Note: Counts of bills by original GLC coding, with missing observations corrected according to RA scoring.
S.4 Measuring Firearm-Related Policy Preference
@@
As for survey responses from state legislators on firearms policy, the most commonly used source is the National Rifle Association (NRA) scorecards, which rate state legislators’ favorability towards firearms policies from self-reported surveys. [^22] The NRA sent surveys to candidates to assess their alignment with NRA core values, and grade them from a F to A+ letter grade scale. Upon close examination of the NRA scores, they suffer from severe missingness and inconsistencies in measurement that pose issues that ultimately prevent us from analyzing them. [^23] See, for example, https://www.thetrace.org/2016/11/nra-gun-record-rating-system-straight-a-students. However, as Figure S.4.1 shows, the Gun Control Score tracks more variation in legislators’ scoring, as most legislators’ NRA ratings stayed unchanged from past elections. At the same time, when a legislator’s NRA score changes, we find that the Gun Control Score also moves in the same direction.
Figure S.4.1: NRA Scoring Compared to Gun Control Score
@@
Note: This figure plots the correlation between the NRA scoring for legislators in California (where available) and the Gun Control Score. We convert an A+ rating to a score of 100, an F rating to 0, and scale the intermediate scores proportionally.
S.5 Estimation Strategy for California Legislators
Legislative sessions in California start on the first Monday in December following an election to convene for an organizational meeting, though the chambers thereafter recess until the first Monday in January, whereupon the legislature is in session year-round until November 30th of the following even-numbered year. Considering each legislative session therefore spans two calendar years, we calculated annual Gun Control Scores for each legislator in our study, so that we could estimate the more immediate effects of mass shootings on voting behavior by considering a legislator’s behavior over the course of one year rather than two. [^24] In analyses where we expand the number of states included in our estimations, we adopt a similar strategy of splitting sessions by years for Colorado, Florida, and Missouri. Texas and Georgia, on the other hand, meet once every two years, so we only have Gun Control Scores for odd-numbered calendar years. Having more votes within a year (and more votes spread throughout the year) would allow for a more precise identification strategy for treatment timing (i.e., monthly effects rather than annual), but this also serves to generally account for the relatively slow pace at which legislation is introduced and deliberated. However, one issue we cannot account for is the concern that votes in the first year of a session may be more procedural than votes in the later year.
S.5.1 Parallel Trends Assumption Holds Overall
Figure S.5.1: Parallel Trends Assumption Holds for California
@@
Figure S.5.2: Sensitivity Analysis of the Parallel Trends Assumption in California Using Smoothness Restrictions
@@
Figure S.5.3: Sensitivity Analysis of the Parallel Trends Assumption in California Using Relative Magnitudes Restrictions
@@
S.5.2 Alternative Estimation Strategy
[…]
where $y_{it}$ denotes the Gun Control Score for legislator $i$ in year $t$ — the average treatment effect on the treated units (ATT). Under this specification, we utilize relative-time indicators (i.e., “Years Since Shooting”) for treatment ($D_{it}^{l}$) instead of a binary treatment indicator (i.e., $D_{it}$) from a canonical TWFE DiD regression. Here, $D_{it}^{l}$ is an indicator for legislator i in group $E_{i}$ being k periods from the start of treatment (i.e., $D_{it}^{l}=\mathbb{I}[t-E_{i}=k]$), such that the first summation captures time periods leading to treatment, and the second summation captures those following treatment.
[…]
S.5.3 Shootings Occurring Within a Legislator’s District Do Not Affect Their Voting Behavior
Figure S.5.4 presents the results of our stacked DiD estimation following Cengiz et al. (2019), for California legislators. We report the results of three models following the specification outlined in Equation 3: one comparing all treated legislators to all others regardless of party, and two that compare treated and untreated legislators within party. Across the specifications for all legislators, Democrats, and Republicans, we find null results — this suggests that the occurrence of a mass shooting within a legislator’s district does not cause a legislator to change their voting behavior on firearm-related policies. Relative to all other legislators, as well as legislators within one’s party, legislators who experience a shooting in their district do not change their voting behavior.
We limit the time periods relative to treatment that we report to those that contain at least one treated legislator in each model specification, and report the formal estimates in Table S.5.1. Due to California’s relatively strict term limit law (12 years in any legislative chamber), we are ultimately limited in assessing the long-running effects of a shooting as legislators tend to be term-limited from holding office before reaching, say, $t=10$ and having pre-treatment observations necessary for our analysis.
Figure S.5.4: ATT Estimates for Stacked DiD Demonstrate Null Results Across Party
Note: We report the null effects of a mass shooting on the voting behavior of CA state legislators on firearm-related policy, with models fit on all legislators, only Democratic legislators, and only Republican legislators. 95% confidence intervals are based on heteroskedasticity-robust standard errors clustered at the legislator and year level. The period immediately prior to treatment, $k=-1$, is excluded to account for multicollinearity.
[…]
In sum, this finding can be interpreted as evidence that (1) mass shootings do not affect the Gun Control Scores of legislators who represent districts in which shootings occur compared to all other legislators; (2) mass shootings do not affect the Gun Control Scores of Democrats representing affected districts relative to other Democrats; and (3) mass shootings do not affect the Gun Control Scores of Republicans representing affected districts relative to other Republicans. However, due to some of the estimation limitations and issues with bias in the ATT that the Callaway and Sant’Anna (2021) estimator resolves, we turn to this alternative estimation strategy to validate the robustness of our finding before making a final conclusion about the effects of a mass shooting on a legislator’s Gun Control Score for those representing the affected districts.
Table S.5.1: ATT Estimates for Stacked and Staggered DiD in California
Overall Democrats Republicans
Time (1) (2) (3) (4) (5) (6)
-4 -0.093 0.045 -0.080 0.167 -0.022 -0.077
(0.056) (0.067) (0.049) (0.049) (0.108) (0.017)
-3 -0.014 0.017 -0.161 -0.131 0.018 0.018
(0.080) (0.058) (0.070) (0.069) (0.046) (0.084)
-2 0.088 0.093 0.013 0.127 0.072 0.047
(0.033) (0.068) (0.043) (0.102) (0.039) (0.035)
-1 -0.062 0.009 -0.030
(0.082) (0.033) (0.039)
0 -0.019 -0.029 -0.087 -0.096 0.024 0.018
(0.057) (0.062) (0.079) (0.079) (0.035) (0.032)
1 -0.057 0.055 -0.065 -0.057 0.020 -0.002
(0.066) (0.106) (0.080) (0.088) (0.076) (0.080)
2 -0.058 0.025 -0.078 -0.075 -0.089 -0.127
(0.058) (0.084) (0.058) (0.054) (0.082) (0.091)
3 -0.098 -0.022 -0.007 0.019 -0.041 -0.120
(0.068) (0.109) (0.067) (0.053) (0.071) (0.067)
4 -0.057 0.071 0.123 0.141 -0.120 -0.204
(0.079) (0.140) (0.057) (0.079) (0.098) (0.091)
Agg. ATT -0.122 0.020 -0.145
(0.072) (0.037) (0.068)
N 8,590 255 6,174 172 2,416 83
Leg. x Chamber FE
Year FE
Est. Strategy CDLZ CS CDLZ CS CDLZ CS
  • Note: Table records the estimated effects of mass shootings occurring within a legislator’s district on a legislator’s Gun Control Score for varying time units (with standard errors reported in parentheses). Estimation strategy refers to the difference-in-differences design utilized in each estimation, with CDLZ referring to the stacked DiD following Cengiz et al. (2019), and CS referring to the staggered DiD following Callaway and Sant’Anna (2021). For the estimations using the CDLZ design, the number of observations counts all repeated control units, which inflates the number of observations through the stacking of datasets due to repeated inclusion. CDLZ leaves out time period -1 to account for multicollinearity. For CS, we use group/cohort aggregation to determine the aggregated ATT and report these estimates. Due to the aggregation by group and time of treatment in CS, year-fixed effects are superfluous, though we include these fixed effects in the CDLZ models.
S.5.4 How Stable is the Gun Control Score for Treated Legislators?
Figure S.5.5 plots the annual Gun Control Score for only the legislators who had a shooting occur within their district, subset by party. The large dots represent averages of the Gun Control Score binned by year and party, with the underlying individual scores being denoted by the x points.
Figure S.5.5: Average Gun Control Score for California Legislators Treated Within District over Time, by Party
Note: Large points are the average scores for legislators of a given party at time t, with smaller x points representing individual treated legislators. Linear regressions have been fit to the averages, not the underlying points, and are subset by pre- and post-treatment.
S.5.5 Results Using IRT Estimates
[…]
Figure S.5.6: Bayesian IRT Estimates Compared to Gun Control Score

### v2

@@ 7 Do Legislators Respond to Shootings in Other Ways?
In the California State Legislature, legislators frequently sign bills as a cosponsor as a means of signaling support for a bill in a low-cost manner. We manually collected the sponsor and cosponsor lists of all firearm-related bills in both chambers of the California State Legislature from 2011 to 2022. Following Fowler (2006), we construct a directed cosponsorship network, with each edge drawn from a cosponsor of a bill to its sponsor. Figure A.10.1 describes this network. This network is densely interconnected, particularly among Democrats, suggesting partisan homophily in the sponsorship network. Table A.10.1 records the most “important” legislators in each session and chamber, measured by highest Outdegree (the number of bills a given legislator cosponsors in a year) centrality, Betweenness centrality, and Eigenvector centrality.
How might mass shootings affect the sponsorship and cosponsorship behaviors of California lawmakers? Beyond Outdegree (which only captures the cosponsorship activities), we define “Activity” as the fraction of firearm-related bills each legislator sponsors and cosponsors out of all firearm-related bills in a year. Together, these measures serve as proxies for a legislator’s level of activity in the gun policy domain, along distinct albeit complementary dimensions. In Table A.10.2, we report the results of our stacked DiD on these two outcomes of interest — Activity and Outdegree. We only find a statistically significant decrease in both the overall Activity measure and the Outdegree measure for treated Democrats, relative to other Democrats, following a shooting in the first post-treatment period. This suggests that treated Democratic legislators in California are less active in the gun policy domain, on average, in the year immediately following a mass shooting. This estimate, however, may be biased by possible pre-trending observed in the pre-treatment periods, suggesting a violation of the parallel trends assumption.
Of course, it is possible that legislators engage in other forms of behavior outside of the legislative process that we may fail to capture in this paper. In subsequent work, we aim to consider the effects of mass shootings on the public-facing behavior of lawmakers — such as social media activity and visits to or meetings with affected constituencies — rather than just the more “official” forms of behavior our current study focuses on.
@@ 8 Conclusion
In California, we find a statistically insignificant 1.15% reduction in treated legislator’s support for restrictive firearm-related policies, compared to all other legislators, an insignificant 4.7% reduction among California Democrats, and an insignificant 1.1% increase among California Republicans. This null finding persists when we expand our definition of treatment to consider the role of geographic proximity and the effects of mass shootings on the voting behavior of legislators who represent districts near the shooting’s location and not just the district in which the shooting occurred. When we expand our analysis to include Colorado, Florida, Missouri, and Texas, we again estimate statistically insignificant effects of mass shootings on legislators compared to members of their own party. The statistically significant results in Georgia are in line with the results we report in Section 6 — the aggregation method underlying the stacked DiD design masks the heterogeneity of treatment effects of individual shootings and our disaggregated results suggest that legislators’ responses vary markedly by incident. Our scoring regime ultimately covers a total of 1,864 unique, individual legislators across six states from 2011 to 2022.
@@
Our findings challenge canonical theories that posit focusing events, such as mass shootings, as critical junctures for changes in legislative behavior. Instead, the results highlight the limits of these events in altering entrenched partisan dynamics, suggesting that focusing events may have differential impacts based on the issue domain and political context. Legislators we interviewed told us that mass shootings led them to sponsor additional gun safety legislation, or reconsider how they could best represent the victims of mass shootings, but also admitted that pressures from party leaders constrained their autonomy in responding to the tragedy at hand. Others suggested that political differences within their constituencies led them to walk a “fine line” as a representative — they could not push a policy agenda at odds with the wishes of the families of victims, for instance. This was in stark contrast to others who told us that the tragedy offered an opportunity to push through policy changes deemed “great” by some legislators and “unconstitutional” Trojan horses by others. Heterogeneity persisted across our quantitative and qualitative results.
@@
This nuance helps contextualize our finding that individual shootings produce differential effects. Some scholarship suggests that particular mass shootings may lead to unique responses, with characteristics such as the race of victims leading to disparate legislative reactions or swings in public opinion (Markarian, 2024; Walker, Collingwood and Bunyasi, 2020). Future studies focusing on other policy domains could assess whether the limited effects estimated in this study are unique to firearm-related issues or generalizable to other policy areas marked by polarization. In this sense, policy-specific ideal points estimates need not be tailored to firearms-related policies alone.
@@
Partisan polarization has contributed to a stagnant firearms policy environment at the federal level, and our findings suggest that this may persist at the sub-national level both within the California Legislature and across five other states with varying gun policy environments. Our results suggest that while state legislatures are active on gun policy issues, mass shootings do not cause any measurable changes in individual-level voting behaviors of elected officials, on average. Every legislator we spoke with said, at different points, that they felt a degree of responsibility to act in the aftermath of the shooting, and yet our results suggest that these events, on average, do not cause a legislator to change their positions on firearm-related policies. However, our disaggregated results could also be interpreted as evidence that individual legislators are responsive to mass shootings, even as aggregating across shootings and legislators washes this effect away. Further, our findings do not indicate that a state’s firearm policy environment is necessarily stagnant — shootings may be changing the scope of legislation that lawmakers vote on (Luca, Malhotra and Poliquin, 2020), but that is beyond the scope of our current study.
@@
This has mixed implications for the state of representative democracy, broadly construed. On the one hand, the stagnation we observe in the voting behavior of state legislators suggests that they are not accountable, on average, to focusing events that may otherwise be thought to incite changes in their behaviors. On the other hand, perhaps this is to be expected, given the relative stability of public opinion on firearm policies after mass shootings. Regardless, while we do not intend to claim that individual votes on niche policies by state legislators capture the attention of the mass public, the outcomes of these votes do have long-lasting and far-reaching implications for accountability and representation.
@@
A.1 Qualitative Interview Procedures
@@
A.2 Locations and Details of Mass Shootings
In this paper, we consider the effect of mass shootings on legislative behavior in California from the 2011-2012 legislative session to the 2021-2022 session. To that end, we focus on shootings occurring between 2011 and 2022 in the state, though we also include the 2010 North Hollywood shooting that may have reasonably affected legislators serving in the 2011-2012 session. As these legislators would reasonably be considered treated in these early sessions, we cannot use them as control units in our later estimations. Table A.2.1 identifies the shootings that are included in our study — specifying the date, location, number of victims (not including the shooter), and the State Senate and Assembly districts in which the shooting occurred (accounting for redistricting following the 2010 Census).
Table A.2.1: Shootings Included in Study and Districts Immediately Affected
@@ A.2 Locations and Details of Mass Shootings
  • Note: Data on shooting location and number of victims (not including the shooter) comes from the Violence Project. Upper refers to the upper chamber district in which the shooting occurred and Lower refers to the respective lower chamber district. Both of these districts are identified through mapping the latitude/longitude location of a shooting to the appropriate state legislative district shapefiles from the United States Census Bureau. These locations are cross-referenced with records from the Gun Violence Archive, when available. The 2010 North Hollywood, CA shooting is effectively dropped from our analyses – as we do not have pre-shooting observations, we cannot consider the effects of the shooting on the voting behavior of affected legislators, so the legislators representing this district that appear in our data from 2011-2022 are considered as previously treated, and removed from subsequent models. Post-Treat reports the year that is considered as $k=0$ in our formal models, and is assigned as the year of the shooting date if the date of the shooting occurs within the first six months of the year, or the subsequent year following the year of the shooting date if the shooting occurs within the last six months of the year. For shootings occurring before the redistricting done following the 2010 Census, we report the district that was affected under the earlier maps, and cross-reference district boundaries after the Census to ensure that legislators who were redistricted to represent a new area are still considered treated if they previously represented an affected district. District shapefiles from this earlier period are sourced primarily from the Redistricting Data Hub.
@@
A.3 Validating the GLC Legislative Scoring Regime
@@
A.4 Measuring Firearm-Related Policy Preference
@@ A.4.1 NRA Ratings
As for survey responses from state legislators on firearms policy, the most commonly used source is the National Rifle Association (NRA) scorecards, which rate state legislators’ favorability towards firearms policies from self-reported surveys. [^20] The NRA sent surveys to candidates to assess their alignment with NRA core values, and grade them from a F to A+ letter grade scale. Upon close examination of the NRA scores, they suffer from severe missingness and inconsistencies in measurement that pose issues that ultimately prevent us from analyzing them. [^21] See, for example, https://www.thetrace.org/2016/11/nra-gun-record-rating-system-straight-a-students. However, as Figure A.4.1 shows, the Gun Control Score tracks more variation in legislators’ scoring, as most legislators’ NRA ratings stayed unchanged from past elections. At the same time, when a legislator’s NRA score changes, we find that the Gun Control Score also moves in the same direction.
Figure A.4.1: NRA Scoring Compared to Gun Control Score
@@
Note: This figure plots the correlation between the NRA scoring for legislators in California (where available) and the Gun Control Score. We convert an A+ rating to a score of 100, an F rating to 0, and scale the intermediate scores proportionally. The correlation is -0.84.
A.4.2 Bayesian IRT Scores
[…]
Figure A.4.2: Bayesian IRT Estimates Compared to Gun Control Score
A.5 Results Using Staggered Difference-in-Differences
A.5.1 Staggered Difference-in-Differences
$ATT(g,t)=\mathbb{E}[Y_{t}(g)-Y_{t}(0)|G_{g}=1]$ (3)
where $g$ denotes a particular group of treated units at time period $t$. Hence, $ATT(g,t)$ is the expected difference between the observed Gun Control Score for treated legislators and untreated legislators at time t. In other words, this yields the group-time average treatment effect (Callaway and Sant’Anna, 2021; Baker, Larcker and Wang, 2022).
where $y_{it}$ denotes the Gun Control Score for legislator $i$ in time $t$ and group $g$. $\alpha_{2}^{g,t}$ represents the effect associated with group g at time t, with $\mathbb{I}[E_{i}=g]$ indicating if legislator i belongs to group g. We also include a time indicator effect, $\alpha_{3}^{g,t}\times\mathbb{I}[T=t]$, where $\alpha_{3}^{g,t}$ represents the effect associated with time t for group g. Using never-treated legislators as our control units, $\beta^{g,t}$ serves as an estimator for $ATT(g,t)$, as this term represents the interaction effect for group g at time t. Ultimately, the unbalanced nature of our panel poses some issues for estimation insofar as we are limited in our ability to explicitly model the outcome evolution for individual legislators who drop out of (and, sometimes, back into) the data. For this reason, we rely on the doubly-robust approach, which improves the robustness of our model against possible misspecification and has no impact on our identification (Callaway and Sant’Anna, 2021; Sant’Anna and Zhao, 2020).
A.5.2 Shootings Occurring Within a Legislator’s District Do Not Affect Their Voting Behavior
Figure A.5.1 presents the results of our staggered DiD estimation following (Callaway and Sant’Anna, 2021), for California legislators. We report the results of three models following the specification outlined in Equation 3: one comparing all treated legislators to all others regardless of party, and two that compare treated and untreated legislators within party. Across the specifications for all legislators, Democrats, and Republicans, we find null results — this suggests that the occurrence of a mass shooting within a legislator’s district does not cause a legislator to change their voting behavior on firearm-related policies. Relative to all other legislators, as well as legislators within one’s party, legislators who experience a shooting in their district do not change their voting behavior.
Figure A.5.1: ATT Estimates for Staggered DiD Demonstrate Null Results Across Party
We limit the time periods relative to treatment that we report to those that contain at least one treated legislator in each model specification, and report the formal estimates in Table A.5.1.
[…]
We report our estimates across our two DiD designs in Table A.5.1 and find very similar estimates at t = 0.
In sum, this finding can be interpreted as evidence that (1) mass shootings do not affect the Gun Control Scores of legislators who represent districts in which shootings occur compared to all other legislators; (2) mass shootings do not affect the Gun Control Scores of Democrats representing affected districts relative to other Democrats; and (3) mass shootings do not affect the Gun Control Scores of Republicans representing affected districts relative to other Republicans.
Table A.5.1: ATT Estimates for Stacked and Staggered DiD in California
Overall Democrats Republicans
Time (1) (2) (3) (4) (5) (6)
-4 -0.108 (0.065) 0.048 (0.069) -0.084 (0.056) 0.177 (0.062) -0.01 (0.11) -0.081 (0.021)
-3 -0.04 (0.099) 0.018 (0.063) -0.17 (0.081) -0.139 (0.074) 0.022 (0.034) 0.019 (0.093)
-2 0.067 (0.031) 0.099 (0.071) 0.017 (0.049) 0.136 (0.11) 0.071 (0.026) 0.05 (0.041)
-1 -0.065 (0.088) 0.011 (0.037) -0.031 (0.047)
0 -0.023 (0.066) -0.03 (0.063) -0.094 (0.09) -0.1 (0.086) 0.022 (0.03) 0.019 (0.038)
1 -0.061 (0.073) 0.058 (0.11) -0.072 (0.078) -0.058 (0.095) 0.022 (0.072) -0.002 (0.084)
2 -0.061 (0.056) 0.027 (0.091) -0.086 (0.06) -0.077 (0.061) -0.098 (0.061) -0.135 (0.093)
3 -0.102 (0.08) -0.022 (0.116) -0.011 (0.086) 0.022 (0.063) -0.044 (0.07) -0.127 (0.076)
4 -0.045 (0.092) 0.076 (0.15) 0.13 (0.054) 0.153 (0.083) -0.129 (0.096) -0.216 (0.099)
# of Legislators 234 255 156 172 78 83
Legislator FE
Chamber FE
Year FE
Est. Strategy CDLZ CS CDLZ CS CDLZ CS
  • Note: Table records the estimated effects of mass shootings occurring within a legislator’s district on a legislator’s Gun Control Score for varying time units (with standard errors reported in parentheses). Estimation strategy refers to the difference-in-differences design utilized in each estimation, with CDLZ referring to the stacked DiD following Cengiz et al. (2019), and CS referring to the staggered DiD following Callaway and Sant’Anna (2021). CDLZ leaves out time period -1 to account for multicollinearity. Due to the aggregation by group and time of treatment in CS, year-fixed effects are superfluous, though we include these fixed effects in the CDLZ models. We cluster standard errors at the legislator level for CS models.
A.5.3 Parallel Trends Assumption Holds Overall
Figure A.5.2: The Average Effect by Length Of Exposure for California
@@ A.5.3 Parallel Trends Assumption Holds Overall
Figure A.5.3: Sensitivity Analysis of the Parallel Trends Assumption in California Using Smoothness Restrictions
@@
Figure A.5.4: Sensitivity Analysis of the Parallel Trends Assumption in California Using Relative Magnitudes Restrictions
@@
A.6 Results Using Alternative Measurement Strategies
A.6.1 Results Using IRT Estimates